1) Use your data to calculate the initial ball velocity for each of the HORIZONTAL launcher settings: Part 2 - Initial Ball Velocity Vertical Distance Small Range Trial 1 Total Horizontal Distance (cm) 161.4 Launch Velocities (
Answer #1

como es un lanzamiento horizontal, la velocidad inicial es solo horizontal, que es constante en todo el movimiento y solo varía en el eje-Y. Por lo tanto, puede calcular el valor inicial por.

x=v_{0x}t_f

clearing the variable initial velocity, we have

v_{0x}=\frac{x}{t_f}

then, taking the average values of the distance X and time of flight, the initial velocity for each range is:

1. Small Range

v_{0x}=\frac{x}{t_f}=\frac{1.6087m}{0.453s}=3.551m/s

2. Medium Range

v_{0x}=\frac{x}{t_f}=\frac{1.931m}{0.453s}=4.263m/s

3. Long Range

v_{0x}=\frac{x}{t_f}=\frac{2.185m}{0.453s}=4.823m/s